OpenStudy (lilshane):

hich related equation can be used to solve the equation? Column A Column B 1. 8 + a = 3.9 2. 3.9 + a = 8 3. 3.9 + a = –8 4. 8 + a = –3.9 A. a = –3.9 – 8 B. a = 3.9 – 8 C. a = 8 – 3.9

OpenStudy (lilshane):

@sierra!

OpenStudy (sierra!):

im not quite sure

OpenStudy (sierra!):

@Directrix can you help him please im not sure about this

OpenStudy (lilshane):

@Directrix

Directrix (directrix):

Sorry, but I don't understand the question. I do not know what is in column A and column B. Do you think you could write this as it appears in your text? You could click on the blue DRAW tab and write it by hand. @lilshane

OpenStudy (lilshane):

the first 4 are the 1 row th elast three are the 2 row

OpenStudy (lilshane):

@Directrix

Directrix (directrix):

Which related equation can be used to solve the equation? 1. 8 + a = 3.9 A. a = –3.9 – 8 2. 3.9 + a = 8 B. a = 3.9 – 8 3. 3.9 + a = –8 C. a = 8 – 3.9 4. 8 + a = –3.9 @lilshane Is this the problem and is it a matching problem?

OpenStudy (lilshane):

How did the Euphrates River contribute to the development and splendor of Babylon? A. The cedar forests on its banks supplied wood for the construction. B. The gold and silver it supplied financed the construction. C. It brought trade, irrigation water, and building materials. D. It protected Babylon from invaders. @Directrix